OpenStudy (anonymous):

The denominator of a fractor is greater than the numerator by 3. If 7 is added to the numerator, and 5 is added to the denominator, then thee fraction itself is increased by 1/2. Find the original fraction.

OpenStudy (anonymous):

@amistre64 @saifoo.khan @sammmmm @ganeshie8

OpenStudy (amistre64):

\[\frac nd~:~n=d+3\] \[\frac {d+3}d\] \[\frac {d+3+7}{d+5}=\frac12\]

OpenStudy (amistre64):

solvinf for d is simple enough to do

OpenStudy (amistre64):

might have a slight error at the end :)

OpenStudy (amistre64):

\[\frac {d+3}d=k\]\[\frac {d+3+7}{d+5}=\frac k2\]

OpenStudy (amistre64):

or is that k + 1/2?

OpenStudy (shubhamsrg):

i think it means d = n+3 and yes it should be k+ 1/2

OpenStudy (amistre64):

lol, i agree :)

OpenStudy (anonymous):

i cant find what d is equall to